A fair pair of dice is rolled

36

times. Find the probability that the same point comes up on both dice at least

11

times
Answer #1

P(a getting a same point onboth dice) = 6/36 = 1/6

n = 36

mean = n * p = 36 * 1/6 = 6

sd = sqrt(n * p * (1 - p))

      = sqrt(36 * 1/6 * 5/6)

      = 2.236

P(X > 11)

= P((X - mean)/sd > (11 - mean)/sd)

= P(Z > (11 - 6)/2.236)

= P(Z > 2.24)

= 1 - P(Z < 2.24)

= 1 - 0.9875

= 0.0125